Jenny puts aside $20 at the end of each month for 3 years. How much will she have then of the investment earns 8.2% p.a., paid monthly?
Found 2 solutions by Theo, mananth:
Answer by Theo(13342)

the investment is 20 dollars at the end of each month for 3 years = 3 * 12 = 36 months.
the interest rate if 8.2% per month compounded monthly = 8.2% / 12 = .6833333333...% per month.
it's a never ending decimal, but carrying out the 3's for 8 positions is usually sufficient.

using the online calculator at https://arachnoid.com/finance/, i get a future value of $813.16 after i click on FV.

my inputs and outputs are shown below, using this online calculator.

$$$

$$$








Answer by mananth(16946)



FV = PMT * ((1+r)^n - 1))/r,
PMT =20
r= 8.2% = 0.082/12 =0.00683
FV = (20*((1+0.00683)^36 -1))/0.00683
FV = $ 813.11 approx